OpenStudy (anonymous):

if f(x)=5x and g(x)=5x/3 find f(x) divided by g(x)

OpenStudy (anonymous):

\[\frac{ 5x }{ \frac{ 5x }{ 3 } }\] \[ 5x *\frac{ 3 }{ 5x }\] 3

OpenStudy (anonymous):

I'll go with 3, thanks My daughter got 1/3

OpenStudy (anonymous):

here is a general solution you can use to check your daughter homework \[\frac{ a }{ \left( \frac{ b }{ c } \right) } = a * \frac{ c }{ b }\]